OpenStudy (anonymous):

2) Seven times a number is subtracted from 15. The result is 42

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

7x-15=42 7x=42+15=57 x= 57/7

OpenStudy (anonymous):

ahhh i see. haha

OpenStudy (anonymous):

:D

OpenStudy (angela210793):

if it is 7times a number subtracted by 15 then it must be 15-7x=42 -7x=42-15 -7x=27 x=-27/7